OpenStudy (anonymous):

I just need to solve 2sqrt5(105) please

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@peachpi

OpenStudy (astrophysics):

\[2\sqrt{5(105)}?\]

OpenStudy (anonymous):

Yes @Astrophysics

OpenStudy (astrophysics):

Ok so what have you tried?

OpenStudy (anonymous):

I don't know how to solve big square roots really. But I need this now :( @Astrophysics

OpenStudy (astrophysics):

\[2\sqrt{525}\] maybe that's easier, just keep simplifying that

OpenStudy (anonymous):

Okay I got 10 times sqrt21 @Astrophysics

OpenStudy (astrophysics):

Yes! Exactly

OpenStudy (anonymous):

That's the answer ?

OpenStudy (astrophysics):

Yup :)

OpenStudy (anonymous):

Okay well I'm solving this problem LAW ENFORCEMENT:  A police accident investigator can use the formula S25L to estimate the speed s of a car in miles per hour based on the length l in feet of the skid marks it left.  How fast was a car traveling that left skid marks 105 feet long? And I have to round the answer to the nearest mph @ Astrophysics

OpenStudy (anonymous):

@Astrophysics

OpenStudy (astrophysics):

What is S25L

OpenStudy (anonymous):

S=2sqrt5l

OpenStudy (astrophysics):

Oooh \[S = 2\sqrt{5L}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I need this answer in the next 3 minutes :(

OpenStudy (astrophysics):

So S is your speed and L is your length this is exactly the same as your problem above

OpenStudy (astrophysics):

plug L = 105

OpenStudy (astrophysics):

I guess you don't have to convert any units either ft into miles or what ever does look like it

OpenStudy (anonymous):

so that's the answer 10 times sqrt ?

OpenStudy (anonymous):

Sqrt21 ? So there is no rounding to the nearest mph

OpenStudy (astrophysics):

Then put it in decimals if it's asking in that

OpenStudy (anonymous):

What would it be

OpenStudy (astrophysics):

Use a calculator..

OpenStudy (anonymous):

One minute left :(

OpenStudy (astrophysics):

I got 46

OpenStudy (anonymous):

I did too! Thanks fam !

OpenStudy (astrophysics):

Yw